# On the Provable Suboptimality of Momentum SGD in Nonstationary Stochastic Optimization

### v2

@@ 3.3 Minimax Regret Bounds
$\mathcal{R}_{T}^{\pi}(F)\;:=\;\mathbb{E}\!\left[\sum_{t=0}^{T-1}\Big(F_{t+1}(\bm{\theta}_{t})-F_{t+1}(\bm{\theta}_{t+1}^{\star})\Big)\right].$
To obtain nontrivial guarantees, given a sequence of differentiable functions $f_{1},\dots,f_{T}:\Theta\rightarrow\mathbb{R}$, define the $L_{p,q}$ gradient-variational functional of $f=(f_{1},\dots,f_{T})$ as
$\mathrm{GVar}_{p,q}(f):=\begin{cases}\left(\dfrac{1}{T}\sum_{t=1}^{T-1}\left\|\nabla f_{t+1}-\nabla f_{t}\right\|_{p}^{\,q}\right)^{1/q}&1\leq p\leq\infty,\;1\leq q<\infty\\[4.30554pt] \max_{1\leq t\leq T-1}\left\|\nabla f_{t+1}-\nabla f_{t}\right\|_{p}&q=\infty\end{cases}$ (3.9)
where for any measurable function $f:\Theta\rightarrow\mathbb{R}$, we have
$\|h\|_{p}:=\begin{cases}\left(\displaystyle\int_{\mathcal{X}}\|h(x)\|_{2}^{\,p}\,dx\right)^{1/p}&p<\infty\\[4.30554pt] \displaystyle\sup_{x\in\mathcal{X}}\|h(x)\|_{2}&p=\infty.\end{cases}$ (3.10)
We set a budget constraint defined by the function class
$\mathcal{F}_{p,q}(V_{T}):=\left\{f:\mathrm{GVar}_{p,q}(f)\leq V_{T}\right\}.$ (3.11)
This is again in contrast to [BGZ15, CWW19] who place constraints on the function values rather than the gradients. We do this since gradient variation naturally gives rise to distributional shift metrics such as the the Wasserstein distance $\mathcal{W}_{1}(\Pi_{t+1},\Pi_{t})$ via Kantorovich-Rubenstein duality and the Total Variation (TV) distance. We also note that placing a gradient variation budget implies a budget on the minimizer drift by strong convexity. We now establish the following worst-case regret for noisy gradient feedback over $\mathcal{F}_{p,q}(V_{T})$. We defer the proofs to Appendix˜C. We will denote the minimax risk as $\mathfrak{M}_{T}(\Pi_{\beta},\mathcal{V}_{T})~:=~\inf_{\pi\in\Pi_{\beta}}\ \sup_{F:\,\mathrm{GVar}_{p,q}(F)\leq\mathcal{V}_{T}}\ \mathcal{R}_{T}^{\pi}(F).$
Theorem 3.7 (Minimax lower bound for strongly-convex function sequences using SGDM).
Fix arbitrary $1\leq p<\infty$ and $1\leq q<\infty$. Suppose ˜2.2 and ˜2.3 hold. Consider the class $\Pi_{\beta}$ of $\mathrm{SGDM}(\beta)$ policies with constant step size $\gamma\leq c_{0}(1-\beta)^{2}/L$. Then there exists a class $\mathcal{F}_{p,q}(V_{T})$ of $\mu$-strongly convex, $L$-smooth function sequences whose gradient-variational functional budget satisfies $\mathrm{GVar}_{p,q}(f)\leq\mathcal{V}_{T}$ such that
@@
$\mathfrak{M}_{T}(\Pi_{\beta},\mathcal{V}_{T})\;\gtrsim\;\max\Bigl\{$ $(1-\beta)^{-2/(\alpha q+2)}\sigma^{4/(\alpha q+2)}\mu^{(\alpha q-2q-2)/(\alpha q+2)}\mathcal{V}_{T}^{2q/(\alpha q+2)}T^{\alpha q/(\alpha q+2)},$ (3.12)
$(1-\beta)^{-2/(\alpha q)}\mu^{(\alpha q-2q-2)/(\alpha q)}L^{2/(\alpha q)}\mathcal{V}_{T}^{2/\alpha}T^{1-2/(\alpha q)}\Bigr\}.$

### v3

@@ E.4 Constructing a $J$-block hard family
$D_{\mathrm{KL}}(P_{u}^{\pi}\,\|\,P_{v}^{\pi})\;=\;\frac{1}{2\sigma^{2}}\sum_{t=1}^{T}\mathbb{E}_{u}\!\left[\big\|\nabla G_{t}^{u}(\bm{\theta}_{t-1})-\nabla G_{t}^{v}(\bm{\theta}_{t-1})\big\|_{2}^{2}\right].$ (201)
Moreover, for the block construction with base losses $\{g_{+},g_{-}\}$, we have the uniform bound
$D_{\mathrm{KL}}(P_{u}^{\pi}\,\|\,P_{v}^{\pi})\;\leq\;C\,\frac{\mu^{2}a^{2}}{\sigma^{2}}\,T,$ (202)
for a universal constant $C>0$ (using $\sup_{\bm{\theta}\in\Theta}\|\nabla g_{+}(\bm{\theta})-\nabla g_{-}(\bm{\theta})\|_{2}\leq C^{\prime}\mu a$ from Lemma˜E.5).
Proof of Lemma˜E.12.
$Y_{t}\,\big|\,\mathcal{F}_{t-1}\;\sim\;\mathcal{N}\big(\nabla G_{t}^{w}(\bm{\theta}_{t-1}),\,\sigma^{2}I_{d}\big).$
Using the chain rule for KL divergence,
$D_{\mathrm{KL}}(P_{u}^{\pi}\|P_{v}^{\pi})=\sum_{t=1}^{T}\mathbb{E}_{u}\!\left[D_{\mathrm{KL}}\!\left(P_{u}^{\pi}(Y_{t}\mid\mathcal{F}_{t-1})\,\big\|\,P_{v}^{\pi}(Y_{t}\mid\mathcal{F}_{t-1})\right)\right].$ (203)
Since the two conditional distributions in (203) are Gaussians with the same covariance $\sigma^{2}I_{d}$, their conditional KL is
Substituting $m_{1}=\nabla G_{t}^{u}(\bm{\theta}_{t-1})$ and $m_{2}=\nabla G_{t}^{v}(\bm{\theta}_{t-1})$ into (203) yields (201). By Lemma˜E.6, we have that $\big\|\nabla g_{+}(\bm{\theta})-\nabla g_{-}(\bm{\theta})\big\|_{2}^{2}\;\leq\;4C\mu^{2}a^{2}\mathbf{1}\{\|\bm{\theta}\|_{2}<r\}$. This results in
Combining this with Lemma˜E.8, we conclude that
E.5 Obtaining a minimax lower bound in the statistical/variation-limited regime
Corollary E.2 (Fano lower bound via KL control).
$\log|\mathcal{U}|\ \geq\ \frac{J}{8}\log 2.$ (204)
$\mathbb{P}(\widehat{U}\neq U)\;\geq\;1-\frac{I(U;Z_{1:T})+\log 2}{\log|\mathcal{U}|}.$ (205)
In particular, if the Gaussian-gradient KL bound of Lemma˜E.12 holds so that
$\sup_{u,v\in\mathcal{U}}D_{\mathrm{KL}}(P_{u}^{\pi}\,\|\,P_{v}^{\pi})\;\leq\;C\,\frac{\mu^{2}a^{2}}{\sigma^{2}/(1-\beta)}\,T,$ (206)
$C\,\frac{\mu^{2}a^{2}}{\sigma^{2}/(1-\beta)}\,T\;\leq\;\frac{1}{16}\,J\log 2,$ (207)
then every estimator $\widehat{U}$ obeys the constant error lower bound
$\mathbb{P}(\widehat{U}\neq U)\ \geq\ \frac{1}{2}.$ (208)
Proof of Corollary˜E.2.
Combining Lemma˜G.7 with (206) yields
Substituting this into (205) and using (204) gives
Under the condition (207), the numerator is at most $(J/16)\log 2+\log 2$, hence for all $J\geq 16$ the right-hand side is at least $1/2$ after adjusting constants (and for bounded $J$ this only changes constants in the final lower bound). This yields (208). ∎
We can finally conclude with Lemma˜E.1 and the upper and lower bounds we have on the KL divergence and discrepancy measure $\chi$ respectively.
Lemma E.13 (Parameter tuning for the information-limited lower bound).
Fix $1\leq p\leq\infty$, $1\leq q\leq\infty$, and set $\alpha:=1+d/p$. Consider the $J$-block construction $\{G^{u}_{1:T}:u\in\mathcal{U}\}$ with block length $\Delta_{T}:=\lfloor T/J\rfloor$ and base losses $\{g_{+},g_{-}\}$. Assume:
  1. 1.
    Gradient-variation bound: from Lemma˜E.10, we have $\mathrm{GVar}_{p,q}(G^{u}_{1:T})\ \leq\ C\,\mu\,a^{\alpha}\,J^{1/q},\;\;\;\forall u\in\mathcal{U}$;
  2. 2.
    Gaussian-gradient KL bound: from Lemma˜E.12, we have $\sup_{u,v\in\mathcal{U}}D_{\mathrm{KL}}(P_{u}^{\pi}\,\|\,P_{v}^{\pi})\ \leq\ C\mu^{2}a^{2}T/(\sigma^{2}/(1-\beta)),\;\;\;\forall\pi$;
  3. 3.
    Packing size lower bound: from Lemma˜E.9, we have the packing size lower bound $\log|\mathcal{U}|\geq c\,J$ for a universal constant $c>0$ (e.g., from Varshamov–Gilbert or constant-weight codes).
Let $\mathbb{V}_{T}>0$ be the variation budget. Choose
[…]
Then $\mathrm{GVar}_{p,q}(G^{u}_{1:T})\leq\mathbb{V}_{T}$ for all $u\in\mathcal{U}$. Moreover, if $J$ additionally satisfies
$C\,\frac{\mu^{2}a^{2}}{\sigma^{2}/(1-\beta)}\,T\ \leq\ \frac{c}{2}\,J,$ (210)
then Fano’s method yields a constant testing error (hence a constant separation in regret) for the family $\{G^{u}_{1:T}:u\in\mathcal{U}\}$ under any policy. In particular, substituting (209) into (210) shows it suffices to take
[…]
and for this choice one obtains the scaling
@@ E.5 Obtaining a minimax lower bound in the statistical/variation-limited regime
$\inf_{\pi\in\Pi_{\beta}}\sup_{G:\,\mathrm{GVar}_{p,q}(G)\leq\mathbb{V}_{T}}\mathcal{R}_{T}^{\pi}(G)\;\gtrsim\;(1-\beta)^{-2/(\alpha q+2)}\sigma^{4/(\alpha q+2)}\mu^{(\alpha q-2q-2)/(\alpha q+2)}\mathbb{V}_{T}^{2q/(\alpha q+2)}T^{\alpha q/(\alpha q+2)},$ (212)